# Assessing treatment effects in observational data with missing confounders: A comparative study of practical doubly-robust and traditional missing data methods

### v1

@@ 5 Discussion
Supplemental Figure S24: Synthetic Data MNAR Simulation: Census cOR. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 for the simple outcome models and 0.404 and 0.371 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S25: Synthetic Data MNAR Simulation: Oracle cOR. Comparing estimators of the oracle truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true cOR value is 0.405 in all scenarios.
@@
Supplemental Figure S26: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.037 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S27: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.031 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S28: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.315 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S29: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.271 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S30: Synthetic Data MNAR Simulation: Census cOR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True clogOR values are 0.405 for the simple outcome models and 0.402 and 0.38 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S31: Synthetic Data MNAR Simulation: Oracle cOR. Comparing estimators of the oracle truth with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true cOR value is 0.405 in all scenarios.
@@
Supplemental Figure S32: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.019 for simple outcome models, and 0.016 and 0.019 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S33: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. The true mRD values are 0.019 for simple outcome models, and 0.016 and 0.015 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S34: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.369 for simple outcome models and 0.374 and 0.281 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S35: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$. True values of mlogRR are 0.374 for simple outcome models and 0.369 and 0.349 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
[…]
The full tree has a maximum depth of 15 nodes. Splitting variables not seen in the first four layers include the Patient Health Questionnaire (PHQ) item 9 score, anxiety, Charlson score, mental health inpatient use (MH IP), anxiety, alcohol use, sex, and prior self-harm. Assignment likelihoods range from 9.7% to 91%.
@@
[…]
(a) The full tree for the 5-year outcome has a maximum depth of 13 nodes. Splitting variables not seen in the first four layers include prior self-harm and alcohol use. Outcome likelihoods range from 2.6% to 40.1%. (b) The full tree for the 1-year outcome has a maximum depth of 14 nodes. Splitting variables not seen in the first four layers include Charlson comorbidity score, alcohol use disorder, PHQ-8 total score, and treatment type. Outcome likelihoods range from 15.3% to 0.9%.
@@
[…]
The full tree has a maximum depth of 16 nodes. Splitting variables not seen in the first four layers include mental health inpatient use (MH IP), Charlson score, alcohol use, and prior self-harm. Missingness likelihoods range from 21% to 75%.
@@
Supplemental Figure S39: Plasmode Simulation: Census clogOR. Comparing estimators of the census truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of clogOR from lightest to darkest scenarios are 0.113, 0.017, -0.192, and -0.069.
Supplemental Figure S40: Plasmode Simulation: Oracle clogOR. Comparing estimators of the oracle truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of clogOR from lightest to darkest scenarios are 0.104 and -0.206.
@@
Supplemental Figure S41: Plasmode Simulation: Census mlogRR. Comparing estimators of the census truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of mlogRR from lightest to darkest scenarios are 0.107, 0.016, -0.159, and -0.058.
Supplemental Figure S42: Plasmode Simulation: Oracle mlogRR. Comparing estimators of the oracle truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at .$95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$. True values of mlogRR from lightest to darkest scenarios are 0.106, -0.033, -0.170, and -0.02.

### v2

@@ 5 Discussion
Supplemental Figure S24: Synthetic Data MNAR Simulation: Census cOR. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 for the simple outcome models and 0.404 and 0.371 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S25: Synthetic Data MNAR Simulation: Oracle cOR. Comparing estimators of the oracle truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true cOR value is 0.405 in all scenarios.
@@
Supplemental Figure S26: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.037 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S27: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.040 for simple outcome models, and 0.038 and 0.031 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S28: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.315 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S29: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 12% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.334 for simple outcome models and 0.339 and 0.271 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S30: Synthetic Data MNAR Simulation: Census cOR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True clogOR values are 0.405 for the simple outcome models and 0.402 and 0.38 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S31: Synthetic Data MNAR Simulation: Oracle cOR. Comparing estimators of the oracle truth with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true cOR value is 0.405 in all scenarios.
@@
Supplemental Figure S32: Synthetic Data MNAR Simulation: Census mRD. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.019 for simple outcome models, and 0.016 and 0.019 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S33: Synthetic Data MNAR Simulation: Oracle mRD. Comparing estimators of the oracle truth with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ The true mRD values are 0.019 for simple outcome models, and 0.016 and 0.015 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S34: Synthetic Data MNAR Simulation: Census mlogRR. Comparing estimators of the census estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.369 for simple outcome models and 0.374 and 0.281 for complex outcome models for MNAR unobserved and MNAR value, respectively.
Supplemental Figure S35: Synthetic Data MNAR Simulation: Oracle mlogRR. Comparing estimators of the oracle estimand with 80% confounder missingness and 5% outcome proportion. Top graph: %Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{2500}}.$ True values of mlogRR are 0.374 for simple outcome models and 0.369 and 0.349 for complex outcome models for MNAR unobserved and MNAR value, respectively.
@@
Supplemental Figure S39: Plasmode Simulation: Census clogOR. Comparing estimators of the census truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of clogOR from lightest to darkest scenarios are 0.113, 0.017, -0.192, and -0.069.
Supplemental Figure S40: Plasmode Simulation: Oracle clogOR. Comparing estimators of the oracle truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of clogOR from lightest to darkest scenarios are 0.104 and -0.206.
@@
Supplemental Figure S41: Plasmode Simulation: Census mlogRR. Comparing estimators of the census truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of mlogRR from lightest to darkest scenarios are 0.107, 0.016, -0.159, and -0.058.
Supplemental Figure S42: Plasmode Simulation: Oracle mlogRR. Comparing estimators of the oracle truth. Top graph: Bias (median, IQR, min and max of converged simulations); Middle graph: Robust RMSE (rRMSE), using median bias and MAD; Bottom graphs: Nominal and oracle coverage, respectively, with blue confidence bands at $95\pm 1.96\sqrt{\frac{.05\cdot.95}{1000}}.$ True values of mlogRR from lightest to darkest scenarios are 0.106, -0.033, -0.170, and -0.02.